The volume of a given mass ofgas is 300cm cube of 27degree celcius and 700mmHg. What will be it volume of 36degree celcius and 750mmHg

1
Expert's answer
2022-03-04T09:03:03-0500

Temperatures in Kelvins:

27oC = 273 + 27 = 300 K

36oC = 273 + 36 = 309 K


According to the combined gas law,


"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"


Therefore,


"V_2=\frac{P_1V_1T_2}{T_1P_2}=\frac{700\ mmHg\times300\ cm^3\times309\ K}{300\ K\times750\ mmHg}=288\ cm^3"


Answer: 288 cm3